# LS-category and topological complexity of real torus manifolds and Dold manifolds of real torus type

### 12 Jan 2024 v1

@@ 2. LS-category and topological complexity of real torus manifolds
[24, Corollary 4.6] The $j$-th stage generalized real Bott manifold $B_{j}$ of the tower (2) is a small cover over $\prod_{i=1}^{j}\Delta^{n_{i}}$ where $\Delta^{n_{i}}$ is the $n_{i}$-simplex.
@@
[15, Proposition 2.7] Every small cover over a product of simplices is a generalized real Bott manifold.
@@
where $F_{r_{j}}^{j}:=\Delta^{n_{1}}\times\dots\times\Delta^{n_{j-1}}\times F_{r_{j}}^{\Delta_{j}}\times\Delta^{n_{j+1}}\times\dots\times\Delta^{n_{m}}.$ Let $\mathcal{N}_{j}:=\sum_{s=1}^{j}n_{s}$, for $j=1,\ldots,m$. Thus $\mathcal{N}_{1}=n_{1}$ and $\mathcal{N}_{m}=n$. Let us define $\mathcal{N}_{0}:=0$. Note that $|\mathcal{F}(P)|=n_{1}+1+n_{2}+1+\dots+n_{m}+1=n+m$. Define a $\mathbb{Z}_{2}$-characteristic function on $P$ by the following
$\lambda\colon\mathcal{F}(P)\to\mathbb{Z}_{2}^{n},$ (7)
$[\lambda(F^{1}_{1}),\dots,\lambda(F^{1}_{n_{1}+1}),\lambda(F^{2}_{1}),\dots,\lambda(F^{2}_{n_{2}+1})\dots,\lambda(F^{m}_{1}),\dots,\lambda(F^{m}_{n_{m}+1})]$
@@
$[e_{1},\dots,e_{\mathcal{N}_{1}},\bm{\alpha_{1}},e_{\mathcal{N}_{1}+1},\dots,e_{\mathcal{N}_{2}},\bm{\alpha_{2}},\dots,e_{\mathcal{N}_{m-1}+1},\dots,e_{\mathcal{N}_{m}},\bm{\alpha_{m}}].$
@@
$A:=$ $\begin{pmatrix}\bm{\alpha_{1}^{1}}&\bm{\alpha_{2}^{1}}&\dots&\bm{\alpha_{m}^{1}}\\ \bm{\alpha_{1}^{2}}&\bm{\alpha_{2}^{2}}&\dots&\bm{\alpha_{m}^{2}}\\ \vdots&\vdots&\dots&\vdots\\ \bm{\alpha_{1}^{m}}&\bm{\alpha_{2}^{m}}&\dots&\bm{\alpha_{m}^{m}}\\ \end{pmatrix}_{m\times m}=\begin{pmatrix}\bm{\alpha_{1}}&\bm{\alpha_{2}}&\dots&\bm{\alpha_{m}}\end{pmatrix}_{1\times m}.$ (13)
@@
Next, we recall the cohomology ring of the generalized real Bott manifolds from [11]. We denote the facet $F_{r_{j}}^{j}$ by $F_{\mathcal{N}_{j-1}+r_{j}+j-1}$. Consider an indeterminate $x_{i}$ corresponding to the facet $F_{i}$ for each $i=1,2,\dots,n+m$.
Proposition 2.10 ([11]).
The cohomology ring of the generalized real Bott manifold is given by
$H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})\cong\frac{\mathbb{Z}_{2}[x_{1},\dots,x_{n+m}]}{\tilde{I}+\tilde{J}}.$ (14)
The ideal $\tilde{I}$ is given by
@@
On the other hand, using the arguments in [5, Proposition $5.1$], the matrix $A$ is conjugate to a unipotent lower triangular vector matrix of the following form:
@@
[…]
Thus, the ideal $\tilde{J}$ is generated by the coordinates of the following matrix.
$\left(\begin{array}[]{cccccccccccc}1&0&\cdots&0&1&0&\cdots&0&\cdots&\cdots&0&0\\ 0&1&\cdots&0&1&0&\cdots&0&\cdots&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\ddots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&1&0&\cdots&0&\cdots&\cdots&0&0\\ 0&0&\cdots&0&\beta_{11}^{2}&1&\cdots&1&\cdots&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\ddots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&\beta_{1n_{2}}^{2}&0&\cdots&1&\cdots&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\ddots&\ddots&\vdots&\vdots\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\ddots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&\beta_{11}^{m}&0&\cdots&\beta_{21}^{m}&\cdots&\cdots&0&1\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\ddots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&\beta_{1n_{m}}^{m}&0&\cdots&\beta_{2n_{m}}^{m}&\cdots&\cdots&1&1\\ \end{array}\right)_{n\times(n+m)}\begin{pmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{\mathcal{N}_{1}}\\ y_{1}\\ x_{\mathcal{N}_{1}+1}\\ \vdots\\ y_{2}\\ \vdots\\ \vdots\\ \vdots\\ x_{\mathcal{N}_{m}}\\ y_{m}\\ \end{pmatrix}_{(n+m)\times 1}.$
From the above matrix multiplication, the first $n_{1}$ elements are $x_{1}+y_{1}=0$, $x_{2}+y_{1}=0$, …, $x_{\mathcal{N}_{1}}+y_{1}=0$. Therefore, we get $x_{1}=x_{2}=\cdots=x_{\mathcal{N}_{1}}=y_{1}$. For $j\geq 2$,
$x_{\mathcal{N}_{j-1}+k_{j}}=\beta_{1k_{j}}^{j}y_{1}+\beta_{2k_{j}}^{j}y_{2}+\cdots+\beta_{(j-1)k_{j}}^{j}y_{j-1}+y_{j}$
for $1\leq k_{j}\leq n_{j}$.
@@
Let $M^{n}(P,\lambda)$ be the small cover over the product of simplices $P=\prod_{j=1}^{m}\Delta^{n_{j}}$ and $\tilde{A}$ be the Bott matrix as in (22) corresponding to the characteristic function $\lambda$ on $P$. We recall the cohomology ring of a small cover over a product of simplices from Proposition 2.11.
Now we look at the generators $\Gamma_{j}$’s of the ideal $I$ in the cohomology ring in Proposition 2.11. Note that,
@@
$\Gamma_{j}$ $=x_{\mathcal{N}_{j-1}+1}x_{\mathcal{N}_{j-1}+2}\cdots x_{\mathcal{N}_{j}}y_{j}$
$=(y_{j}+\beta_{11}^{j}y_{1}+\beta_{21}^{j}y_{2}+\cdots+\beta_{(j-1)1}^{j}y_{j-1})(y_{j}+\beta_{12}^{j}y_{1}+\beta_{22}^{j}y_{2}+$
@@
Using the Poincare duality, we have ${x_{1}\cdots x_{\mathcal{N}_{1}}x_{\mathcal{N}_{1}+1}\cdots x_{\mathcal{N}_{2}}\neq 0}$. Now $x_{\mathcal{N}_{1}+1}=\ldots=x_{\mathcal{N}_{2}}$ and all of them are either $y_{2}$ or, $y_{1}+y_{2}$. If all of them are equal to $y_{2}$, then $y_{1}^{n_{1}}y_{2}^{n_{2}}\neq 0$. If all of them are equal to $y_{1}+y_{2}$ then
@@
A similar computation as above implies that $y_{1}^{n_{1}}y_{2}^{n_{2}+1}=0$, since ${x_{1}\cdots x_{\mathcal{N}_{1}}\Gamma_{2}=0}$. Again using Poincare duality ${x_{1}\cdots x_{\mathcal{N}_{1}}x_{\mathcal{N}_{1}+1}\cdots x_{\mathcal{N}_{2}}\cdots x_{\mathcal{N}_{1}}x_{\mathcal{N}_{2}+1}\cdots x_{\mathcal{N}_{3}}\neq 0}$ and we have two other condition $y_{1}^{n_{1}+1}=0$ and $y_{1}^{n_{1}}y_{2}^{n_{2}+1}=0$. Using this
@@
We now use Lemma 2.12 to get the sharp lower bound on the topological complexity of generalized real Bott manifolds.
@@
Let $d_{j}=2^{r_{j}}-1$ such that $n_{j}\leq d_{j}<2n_{j}$ for $j=1,\dots,m$. Let $\mathfrak{a}_{j}:=1\otimes y_{j}-y_{j}\otimes 1$. Then $\mathfrak{a}_{j}$ is in the ideal of the zero-divisors of $H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})$ because $\mathfrak{a}_{j}$ belongs to the kernel of the homomorphism
$\cup\colon H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})\otimes H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})\rightarrow H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2}).$ (23)
Now,
@@
Note that the binomial coefficients $\binom{2^{r_{j}}-1}{i}$ are odd for all $0\leq i\leq d_{j}$. The binomial expansion of $\mathfrak{a}_{j}^{d_{j}}$ contains the term $(y_{j}^{d_{j}-n_{j}}\otimes y_{j}^{n_{j}})$ which is non-zero and there is no other same term in the expression of $\mathfrak{a}_{j}^{d_{j}}$. So $\mathfrak{a}_{j}^{d_{j}}$ is nonzero. Therefore $\mathfrak{a}_{1}^{d_{1}}\cdots\mathfrak{a}_{m}^{d_{m}}$ is non-zero in the kernel of the map in (23). Hence zero-divisors-cup-length of $H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})$ is greater than or equal to $d_{1}+\cdots+d_{m}$. Therefore, we have $\mathrm{TC}(M^{n}(P,\lambda))\geq d_{1}+\cdots+d_{m}+1$, i.e.,
@@
Theorem 2.13 gives a better lower bound of the LS-category of a generalized real Bott manifold if each $n_{j}$ is a power of $2$.
@@
If $n_{j}=2^{r_{j}-1}$ then $r_{j}$ satisfies $n_{j}\leq 2^{r_{j}}-1<2n_{j}$. Using Theorem 2.13, we get
@@
Since $n=n_{1}+\dots+n_{m}$, we have $\mathrm{TC}(M^{n}(P,\lambda))\geq 2n-m+1.$
We now compute the sharp upper bound on $\mathrm{TC}(M^{n}(P,\lambda))$.
@@
The real moment angel manifold corresponding to the polytope $P=\prod_{j=1}^{m}\Delta^{n_{j}}$ is $\prod_{j=1}^{m}S^{n_{j}}$. Let $m_{j}=\frac{n_{j}+1}{2}$. Then, for each $j\in\{i_{1},\dots,i_{k}\}\subset\{1,\ldots,m\}$, there is a free $S^{1}$-action on
@@
defined by $\eta(z_{1},\ldots,z_{m_{j}})\mapsto(\eta z_{1},\ldots,\eta z_{m_{j}})$. Thus, there exists a free $(S^{1})^{k}$-action on $\prod_{j=1}^{m}S^{n_{j}}$. Now $M^{n}(P,\lambda)$ can be realized as the orbit space of the moment angle manifold $\prod_{j=1}^{m}S^{n_{j}}$ by a free $(\mathbb{Z}_{2})^{m}$-action, see [15, Remark 2.3]. Since both of these actions are a multiplication by a complex number, they commute and induce a free action of $(S^{1}/\left<e^{\pi\sqrt{-1}}\right>)^{k}$-action on $M^{n}(P,\lambda)$ if $n_{j}>1$ is odd for each $j\in\{i_{1},\dots,i_{k}\}\subset\{1,\ldots,m\}$. Then, the result follows from [21, Corollary 5.3], since $S^{1}/\left<e^{\pi\sqrt{-1}}\right>\cong S^{1}$. ∎
@@
$12^{r_{1}}+\dots+2^{r_{m}}-m+1\leq\mathrm{TC}(M^{n}(P,\lambda))\leq 2n-k+$ 1 .
@@
This follows from Theorem 2.13 and 2.15. ∎
@@
Let $M^{n}(P,\lambda)$ be a small cover over $P=\prod_{j=1}^{m}\Delta^{n_{j}}$. If $n_{j}=2^{r_{j}}+1$ for all $j\in\{1,2\dots,m\}$ and $r_{j}\geq 1$ then
$12n-3m+1\leq\mathrm{TC}(M^{n}(P,\lambda))\leq 2n-m+$ 1 .
Example 2.18.
  1. (1)
    $47\leq\mathrm{TC}(M^{n}(P,\lambda))\leq 49$ if $P=\Delta^{16}\times\Delta^{8}$.
  2. (2)
    $47\leq\mathrm{TC}(M^{n}(P,\lambda))\leq 50$ if $P=\Delta^{16}\times\Delta^{9}$ or $\Delta^{17}\times\Delta^{8}$.
  3. (3)
    $47\leq\mathrm{TC}(M^{n}(P,\lambda))\leq 51$ if $P=\Delta^{17}\times\Delta^{9}$.

### 04 Apr 2026 v2

@@ 2. LS-category and topological complexity of real torus manifolds
[23, Corollary 4.6] The $j$-th stage generalized real Bott manifold $B_{j}$ of the tower (2) is a small cover over $\prod_{i=1}^{j}\Delta^{n_{i}}$ where $\Delta^{n_{i}}$ is the $n_{i}$-simplex.
@@
[14, Proposition 2.7] Every small cover over a product of simplices is a generalized real Bott manifold.
@@
where $F_{r_{j}}^{j}:=\Delta^{n_{1}}\times\dots\times\Delta^{n_{j-1}}\times F_{r_{j}}^{\Delta_{j}}\times\Delta^{n_{j+1}}\times\dots\times\Delta^{n_{m}}.$ Let $\mathcal{N}_{j}:=\sum_{s=1}^{j}n_{s}$, for $j=1,\ldots,m$. Thus $\mathcal{N}_{1}=n_{1}$ and $\mathcal{N}_{m}=n$. Let us define $\mathcal{N}_{0}:=0$. Note that $|\mathcal{F}(P)|=n_{1}+1+n_{2}+1+\dots+n_{m}+1=n+m$. Define a $\mathbb{Z}_{2}$-characteristic function $\lambda\colon\mathcal{F}(P)\to\mathbb{Z}_{2}^{n},$ on $\mathcal{F}(P)$ by the following.
$\lambda(F^{1}_{1}),\dots,\lambda(F^{1}_{n_{1}+1}),\lambda(F^{2}_{1}),\dots,\lambda(F^{2}_{n_{2}+1})\dots,\lambda(F^{m}_{1}),\dots,\lambda(F^{m}_{n_{m}+1})]$
@@
$e_{1},\dots,e_{\mathcal{N}_{1}},\bm{\alpha_{1}},e_{\mathcal{N}_{1}+1},\dots,e_{\mathcal{N}_{2}},\bm{\alpha_{2}},\dots,e_{\mathcal{N}_{m-1}+1},\dots,e_{\mathcal{N}_{m}},\bm{\alpha_{m}}].$
@@
$A:=$ $\begin{pmatrix}\bm{\alpha_{1}^{1}}&\bm{\alpha_{2}^{1}}&\dots&\bm{\alpha_{m}^{1}}\\ \bm{\alpha_{1}^{2}}&\bm{\alpha_{2}^{2}}&\dots&\bm{\alpha_{m}^{2}}\\ \vdots&\vdots&\dots&\vdots\\ \bm{\alpha_{1}^{m}}&\bm{\alpha_{2}^{m}}&\dots&\bm{\alpha_{m}^{m}}\\ \end{pmatrix}_{m\times m}=\begin{pmatrix}\bm{\alpha_{1}}&\bm{\alpha_{2}}&\dots&\bm{\alpha_{m}}\end{pmatrix}_{1\times m}.$
@@
Next, we recall the cohomology ring of the generalized real Bott manifolds from [10]. We denote the facet $F_{r_{j}}^{j}$ by $F_{\mathcal{N}_{j-1}+r_{j}+j-1}$. Consider an indeterminate $x_{i}$ corresponding to the facet $F_{i}$ for each $i=1,2,\dots,n+m$. The cohomology ring of the generalized real Bott manifold is given by
$H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})\cong\frac{\mathbb{Z}_{2}[x_{1},\dots,x_{n+m}]}{\tilde{I}+\tilde{J}}$, (7)
where the ideal $\tilde{I}$ is given by
@@
On the other hand, using the arguments in [4, Proposition $5.1$], the matrix $A$ is conjugate to a unipotent lower triangular vector matrix of the following form:
@@
where $\bm{\beta_{j}^{k}}=(\beta_{j_{1}}^{k},\beta_{j_{2}}^{k},\dots,\beta_{j_{n_{k}}}^{k})^{t}\in\mathbb{Z}_{2}^{n_{k}}$ and $\textbf{1}=(1,\dots,1)^{t}\in\mathbb{Z}^{n_{k}}_{2}$ for $k=1,\dots,m$. The matrix $\tilde{A}$ is called the Bott matrix. Thus, the ideal $\tilde{J}$ becomes
$\tilde{J}=\big<x_{\mathcal{N}_{j-1}+k_{j}}+\beta_{1k_{j}}^{j}y_{1}+\beta_{2k_{j}}^{j}y_{2}+\cdots+\beta_{(j-1)k_{j}}^{j}y_{j-1}+y_{j}~|~j=1,2,\dots,m\big$>.
In the cohomology ring $H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})$, we get $x_{1}=x_{2}=\cdots=x_{\mathcal{N}_{1}}=y_{1}$. For $j\geq 2$,
$x_{\mathcal{N}_{j-1}+k_{j}}=\beta_{1k_{j}}^{j}y_{1}+\beta_{2k_{j}}^{j}y_{2}+\cdots+\beta_{(j-1)k_{j}}^{j}y_{j-1}+y_{j}\text{ for }1\leq k_{j}\leq n_{j}.$
@@
For the $2$-stage Bott tower over $\Delta^{n_{1}}\times\Delta^{n_{2}}$, in [3, Theorem 4.6] it was shown that $y_{1}^{n_{1}}y_{2}^{n_{2}}\neq 0$. We now generalize this to any $m$-stage Bott tower for arbitrary $m$. This will be used in obtaining a sharp cohomological lower bound on the topological complexity of generalized Bott towers.
@@
Let $M^{n}(P,\lambda)$ be the small cover over the product of simplices $P=\prod_{j=1}^{m}\Delta^{n_{j}}$ and $\tilde{A}$ be the Bott matrix as in (11) corresponding to the characteristic function $\lambda$ on $P$. We recall the cohomology ring of a small cover over a product of simplices from Proposition˜2.10.
Now we look at the generators $\Gamma_{j}$’s of the ideal $I$ in the cohomology ring in Proposition˜2.10. Note that,
@@
$\Gamma_{j}$ $=(y_{j}+\beta_{11}^{j}y_{1}+\beta_{21}^{j}y_{2}+\cdots+\beta_{(j-1)1}^{j}y_{j-1})(y_{j}+\beta_{12}^{j}y_{1}+\beta_{22}^{j}y_{2}+$
@@
Now, using the Poincare duality, we have $x_{1}\cdots x_{\mathcal{N}_{1}}x_{\mathcal{N}_{1}+1}\cdots x_{\mathcal{N}_{2}}\neq 0$}. Now $x_{\mathcal{N}_{1}+1}=\ldots=x_{\mathcal{N}_{2}}$ and all of them are either $y_{2}$ or, $y_{1}+y_{2}$. If all of them are equal to $y_{2}$, then $y_{1}^{n_{1}}y_{2}^{n_{2}}\neq 0$. If all of them are equal to $y_{1}+y_{2}$ then
@@
A similar computation as above implies that $y_{1}^{n_{1}}y_{2}^{n_{2}+1}=0$, since $x_{1}\cdots x_{\mathcal{N}_{1}}\Gamma_{2}=0$}. Again using the Poincare duality $x_{1}\cdots x_{\mathcal{N}_{1}}x_{\mathcal{N}_{1}+1}\cdots x_{\mathcal{N}_{2}}\cdots x_{\mathcal{N}_{1}}x_{\mathcal{N}_{2}+1}\cdots x_{\mathcal{N}_{3}}\neq 0$} and we have two other conditions $y_{1}^{n_{1}+1}=0$ and $y_{1}^{n_{1}}y_{2}^{n_{2}+1}=0$. Using this
@@
We now use Lemma˜2.11 to obtain the sharp lower bound on the topological complexity of generalized real Bott manifolds.
@@
Let $d_{j}=2^{r_{j}}-1$ such that $n_{j}\leq d_{j}<2n_{j}$ for $j=1,\dots,m$. Let $\mathfrak{a}_{j}:=1\otimes y_{j}-y_{j}\otimes 1$. Now,
@@
Note that the binomial coefficients $\binom{2^{r_{j}}-1}{i}$ are odd for all $0\leq i\leq d_{j}$. The binomial expansion of $\mathfrak{a}_{j}^{d_{j}}$ contains the term $(y_{j}^{d_{j}-n_{j}}\otimes y_{j}^{n_{j}})$ which is non-zero and there is no other same term in the expression of $\mathfrak{a}_{j}^{d_{j}}$. So $\mathfrak{a}_{j}^{d_{j}}$ is nonzero. Therefore $\mathfrak{a}_{1}^{d_{1}}\cdots\mathfrak{a}_{m}^{d_{m}}$ is non-zero. Hence zero-divisors-cup-length of $H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})$ is greater than or equal to $d_{1}+\cdots+d_{m}$. Therefore, we have $\mathrm{TC}(M^{n}(P,\lambda))\geq d_{1}+\cdots+d_{m}+1$, i.e.,
@@
Theorem˜2.12 gives a better lower bound of the LS-category of a generalized real Bott manifold if each $n_{j}$ is a power of $2$.
@@
If $n_{j}=2^{r_{j}-1}$ then $r_{j}$ satisfies $n_{j}\leq 2^{r_{j}}-1<2n_{j}$. Using Theorem˜2.12, we get
@@
Since $n=n_{1}+\dots+n_{m}$, we have
$\mathrm{TC}(M^{n}(P,\lambda))\geq 2n-m+1.$
We obtain a sharp upper bound on $\mathrm{TC}(M^{n}(P,\lambda))$, which improves the dimensional upper bound.
@@
The real moment angle manifold corresponding to the polytope $P=\prod_{j=1}^{m}\Delta^{n_{j}}$ is $\prod_{j=1}^{m}S^{n_{j}}$. Let $m_{j}=\frac{n_{j}+1}{2}$. Then, for each $j\in\{i_{1},\dots,i_{k}\}\subset\{1,\ldots,m\}$, there is a free $S^{1}$-action on
@@
defined by $\eta(z_{1},\ldots,z_{m_{j}})\mapsto(\eta z_{1},\ldots,\eta z_{m_{j}})$. Thus, there exists a free $(S^{1})^{k}$-action on $\prod_{j=1}^{m}S^{n_{j}}$. Now $M^{n}(P,\lambda)$ can be realized as the orbit space of the moment angle manifold $\prod_{j=1}^{m}S^{n_{j}}$ by a free $(\mathbb{Z}_{2})^{m}$-action, see [14, Remark 2.3]. Since both of these actions are a multiplication by a complex number, they commute and induce a free action of $(S^{1}/\left<e^{\pi\sqrt{-1}}\right>)^{k}$-action on $M^{n}(P,\lambda)$ if $n_{j}>1$ is odd for each $j\in\{i_{1},\dots,i_{k}\}\subset\{1,\ldots,m\}$. Then, the result follows from [22, Corollary 5.3], since $S^{1}/\left<e^{\pi\sqrt{-1}}\right>\cong S^{1}$. ∎
@@
$2^{r_{1}}+\dots+2^{r_{m}}-m+1\leq\mathrm{TC}(M^{n}(P,\lambda))\leq 2n-k+1.$
@@
This follows from Theorems 2.12 and 2.14. ∎
@@
Let $M^{n}(P,\lambda)$ be a small cover over $P=\prod_{j=1}^{m}\Delta^{n_{j}}$. If $n_{j}=2^{s_{j}}+1$ for all $j=1,\dots,k<m$ and $n_{j}=2^{s_{j}}$ for $j=k+1,\ldots,m$ where each $s_{j}\geq 1$, then
$2n-2k-m+1\leq\mathrm{TC}(M^{n}(P,\lambda))\leq 2n-k+1.$
Proof.
Note that $1\leq n_{j}<2^{s_{j}+1}-1=2n_{j}-3$ for $j=1,\ldots,k$ and $1\leq n_{j}<2^{s_{j}+1}-1=2n_{j}-1$ for $j=k+1,\ldots,m$. Thus, using Theorem˜2.12, we get the lower bound as $\sum_{j=1}^{k}(2n_{j}-3)+\sum_{j=k+1}^{m}(2n_{j}-1)+1=2n-2k-m+1$ . Also, Theorem˜2.14 gives the upper bound. ∎
Remark 2.17.
If $m=3$, the difference between the upper and lower bounds is 3 and 4 when $k=0$ and $k=1$ in Corollary˜2.16, respectively.